### v1

# Constructing an Optimal Behavior Basis for the Option Keyboard

Abstract
Multi-task reinforcement learning aims to quickly identify solutions for new tasks with minimal or no additional interaction with the environment. Generalized Policy Improvement (GPI) addresses this by combining a set of base policies to produce a new one that is at least as good—though not necessarily optimal—as any individual base policy. Optimality can be ensured, particularly in the linear-reward case, via techniques that compute a Convex Coverage Set (CCS). However, these are computationally expensive and do not scale to complex domains. The Option Keyboard (OK) improves upon GPI by producing policies that are at least as good—and often better. It achieves this through a learned meta-policy that dynamically combines base policies. However, its performance critically depends on the choice of base policies. This raises a key question: is there an optimal set of base policies—an optimal behavior basis—that enables zero-shot identification of optimal solutions for any linear tasks? We solve this open problem by introducing a novel method that efficiently constructs such an optimal behavior basis. We show that it significantly reduces the number of base policies needed to ensure optimality in new tasks. We also prove that it is strictly more expressive than a CCS, enabling particular classes of non-linear tasks to be solved optimally. We empirically evaluate our technique in challenging domains and show that it outperforms state-of-the-art approaches, increasingly so as task complexity increases.
1 Introduction
Reinforcement learning (RL) methods have been successfully used to solve complex sequential decision-making problems (Silver et al., , 2017; Bellemare et al., , 2020). However, traditional RL algorithms typically require thousands or millions of interactions with the environment to learn a single policy for a single task. Multi-task RL methods address this limitation by enabling agents to quickly identify solutions for new tasks with minimal or no additional interactions. Such methods often achieve this by learning a set of specialized policies (or behavior basis) designed for specific tasks and subsequently combining them to more rapidly solve novel tasks.
A powerful approach for combining policies to solve new tasks in a zero-shot manner leverages successor features (SFs) and generalized policy improvement (GPI) (Barreto et al., , 2017, 2018, 2020). GPI can combine base policies to solve a new task, producing a policy that is at least as good as any individual base policy. Importantly, however, GPI policies are not guaranteed to be optimal. Optimality can be ensured if tasks can be expressed as a linear combination of reward features using techniques that compute a convex coverage set (CCS; Alegre et al., , 2022). A CCS is a set of policies that includes an optimal policy for any linear task. Unfortunately, the number of policies in a CCS often grows exponentially with the number of reward features (Roijers, , 2016), making existing algorithms computationally expensive and difficult to scale to complex domains (Yang et al., , 2019; Alegre et al., , 2022; Alegre et al., 2023b, ).
The option keyboard (OK) method improves upon GPI by producing policies that are at least as good—and often better. It achieves this through a learned meta-policy that dynamically combines base policies by assigning state-dependent linear weights to reward features (Barreto et al., , 2019). This allows OK to express a larger spectrum of behaviors than GPI and potentially solve tasks beyond linear rewards, enabling it to produce policies that GPI cannot represent. However, its performance critically depends on the choice of base policies available. This raises a key question and open problem: is there an optimal set of base policies for the OK—an optimal behavior basis—that enables zero-shot identification of optimal solutions for any linear tasks? Existing works using OK assume that a set of base policies is either known a priori or constructed with expert domain knowledge and do not address the problem of identifying a good behavior basis (Barreto et al., , 2019; Carvalho et al., 2023b, ).
We solve this open problem by introducing Option Keyboard Basis (OKB), a novel method with strong formal guarantees that efficiently identifies an optimal behavior basis for the OK. Importantly, our method provably identifies a set of policies that allows the OK to optimally solve any linear task; i.e., it can express all policies in a CCS. We show that it significantly reduces the number of base policies required to ensure zero-shot optimality in new tasks. Furthermore, we prove that the set of policies it can express is strictly more expressive than a CCS, enabling particular classes of non-linear tasks to be solved optimally. We empirically evaluate our method in challenging high-dimensional RL problems and show that it consistently outperforms state-of-the-art GPI-based approaches. Importantly, we also observe that the performance gain over competing methods becomes more pronounced as the number of reward features increases.
2 Background
2.1 Reinforcement Learning
An RL problem (Sutton and Barto, , 2018) is typically modeled as a Markov decision process (MDP). An MDP is defined as a tuple $M\triangleq(\mathcal{S},\mathcal{A},p,r,\mu,\gamma)$, where $\mathcal{S}$ is a state space, $\mathcal{A}$ is an action space, $p(\cdot|s,a)$ describes the distribution over next states given that the agent executed action $a$ in state $s$, $r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a reward function, $\mu$ is an initial state distribution, and $\gamma\in[0,1)$ is a discounting factor. Let $S_{t}$, $A_{t}$, and $R_{t}\triangleq r(S_{t},A_{t},S_{t+1})$ be random variables corresponding to the state, action, and reward, respectively, at time step $t$. The goal of an RL agent is to learn a policy $\pi:\mathcal{S}\to\mathcal{A}$ that maximizes the expected discounted sum of rewards (return), $G_{t}=\sum_{i=0}^{\infty}\gamma^{i}R_{t+i}$. The action-value function of a policy $\pi$ is defined as $q^{\pi}(s,a)\triangleq\mathbb{E}_{\pi}[G_{t}|S_{t}=s,A_{t}=a]$, where $\mathbb{E}_{\pi}[\cdot]$ denotes the expectation over trajectories induced by $\pi$. Given $q^{\pi}$, one can define a greedy policy $\pi^{\prime}(s)\in\operatorname*{arg\,max}_{a}q^{\pi}(s,a)$. It is guaranteed that $q^{\pi^{\prime}}(s,a)\geq q^{\pi}(s,a),\forall(s,a)\in\mathcal{S}\times\mathcal{A}$. The processes of computing $q^{\pi}$ and $\pi^{\prime}$ are known, respectively, as the policy evaluation and policy improvement steps. Under certain conditions, repeatedly executing the policy evaluating and improvement steps leads to an optimal policy $\pi^{*}(s)\in\operatorname*{arg\,max}\nolimits_{a}q^{*}(s,a)$ (Puterman, , 2014).
Let $(\mathcal{S},\mathcal{A},p,\mu,\gamma)$ be a Markov control process (MPC; Puterman, , 2014), i.e., an MDP without a reward function. Given an MPC, we define a family of MDPs $\mathcal{M}\triangleq\{(\mathcal{S},\mathcal{A},p,r,\mu,\gamma)\mid r:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}\}$ that only differ in their reward function. We refer to any MDP $M\in\mathcal{M}$ (and its corresponding reward function $r$) as a task.
2.2 Generalized Policy Evaluation and Improvement
Generalized Policy Evaluation (GPE) and Generalized Policy Improvement (GPI) generalize the policy evaluation and improvement steps to the case where an agent has access to a set of policies. In particular, GPE and GPI are used, respectively, (i) to evaluate a policy on multiple tasks; and (ii) to construct a policy (appropriate for solving a particular novel task) by improving upon an existing set of policies.
Definition 2.1 (Barreto et al., (2020)).
GPE is the computation of the action-value function of a policy $\pi$, $q^{\pi}_{r}(s,a)$, on a set of tasks. Given a set of policies $\Pi$ and a reward function of an arbitrary task, $r$, GPI defines a policy, $\pi^{\prime}$, such that
$q^{\pi^{\prime}}_{r}(s,a)\geq\max_{\pi\in\Pi}q^{\pi}_{r}(s,a)\ \text{ for all }(s,a)\in\mathcal{S}\times\mathcal{A}.$ (1)
Based on this definition, for any reward function $r$, a GPI policy can be constructed based on a set of policies $\Pi$ as:
$\pi^{\text{GPI}}(s;\Pi)\in\underset{a\in\mathcal{A}}{\arg\max}\,\,\underset{\pi\in\Pi}{\max}\,\,q^{\pi}_{r}(s,a).$ (2)
Let $q^{\text{GPI}}_{r}(s,a)$ be the action-value function of $\pi^{\text{GPI}}$ under the reward function $r$. The GPI theorem (Barreto et al., , 2017) ensures that $\pi^{\text{GPI}}$ in Eq. (2) satisfies Def. 2.1; i.e., that $q^{\text{GPI}}_{r}(s,a)\geq\max_{\pi\in\Pi}q_{r}^{\pi}(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. This implies that Eq. (2) can be used to define a policy guaranteed to perform at least as well as all other policies $\pi_{i}\in\Pi$ w.r.t. any given reward function, $r$. The GPI theorem can also be extended to the case $q^{\pi_{i}}$ is replaced with an approximation, $\tilde{q}^{\pi_{i}}$ (Barreto et al., , 2018).
2.3 GPE&GPI via Successor Features
Successor features (SFs) allows us to perform GPE&GPI efficiently (Barreto et al., , 2017). Assume that the reward functions of interest are linear w.r.t. reward features, $\bm{\mathbf{\phi}}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{d}$. That is, a reward function $r_{\bm{\mathbf{w}}}$ can be expressed as $r_{\bm{\mathbf{w}}}(s,a,s^{\prime})=\bm{\mathbf{\phi}}(s,a,s^{\prime})\cdot\bm{\mathbf{w}}$, where $\bm{\mathbf{w}}\in\mathbb{R}^{d}$ is a weight vector. Then, the SFs of a policy $\pi$ for a given state-action pair $(s,a)$, $\bm{\mathbf{\psi}}^{\pi}(s,a)\in\mathbb{R}^{d}$, is defined as
$\bm{\mathbf{\psi}}^{\pi}(s,a)\triangleq\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}\gamma^{i}\bm{\mathbf{\phi}}_{t+i}\mid S_{t}=s,A_{t}=a\right],$ (3)
where $\bm{\mathbf{\phi}}_{t}\triangleq\bm{\mathbf{\phi}}(S_{t},A_{t},S_{t+1})$. Notice that the definition of SFs corresponds to a form of action-value function, where the features $\bm{\mathbf{\phi}}_{t}$ play the role of rewards. Thus, SFs can be learned through any temporal-difference (TD) learning algorithm (e.g., Q-learning (Watkins, , 1989; Mnih et al., , 2015)). We write $\bm{\mathbf{\psi}}^{\pi}$ to denote the SF vector associated with $\pi$, where the expectation is with respect to the initial state distribution: $\bm{\mathbf{\psi}}^{\pi}\triangleq\mathbb{E}_{S_{0}\sim\mu}\left[\bm{\mathbf{\psi}}^{\pi}(S_{0},\pi(S_{0}))\right]$.
Given the SFs $\bm{\mathbf{\psi}}^{\pi}(s,a)$ of a policy $\pi$, it is possible to directly compute the action-value function $q^{\pi}_{\bm{\mathbf{w}}}(s,a)$ of $\pi$, under any linearly-expressible reward functions, $r_{\bm{\mathbf{w}}}$, as follows: $q_{\bm{\mathbf{w}}}^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{i=0}^{\infty}\gamma^{i}r_{\bm{\mathbf{w}}}(S_{t+i},A_{t+i},S_{t+i+1})\mid S_{t}=s,A_{t}=a\right]=\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\bm{\mathbf{w}}.$ That is, given any set of reward weights, $\{\bm{\mathbf{w}}_{i}\}_{i=1}^{n}$, GPE can be performed efficiently via inner-products between the SFs and the reward weights: $q_{\bm{\mathbf{w}}_{i}}^{\pi}(s,a)=\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\bm{\mathbf{w}}_{i}$. Note that the value of a policy $\pi$ under any task $\bm{\mathbf{w}}$ can be expressed as $v^{\pi}_{\bm{\mathbf{w}}}=\bm{\mathbf{\psi}}^{\pi}\cdot\bm{\mathbf{w}}$.
Let $\Pi=\{\pi_{i}\}_{i=1}^{n}$ be a set of policies with corresponding SFs $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$. Based on the definition of GPI (Def. 2.1) and the reward decomposition $r_{\bm{\mathbf{w}}}(s,a,s^{\prime})=\bm{\mathbf{\phi}}(s,a,s^{\prime})\cdot\bm{\mathbf{w}}$, a generalized policy, $\pi^{\text{GPI}}:\mathcal{S}\times\mathcal{W}\to\mathcal{A}$, can then be defined as follows:
$\pi^{\text{GPI}}(s,\bm{\mathbf{w}};\Pi)\in\underset{a\in\mathcal{A}}{\arg\max}\,\,\underset{\pi\in\Pi}{\max}\,\,\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\bm{\mathbf{w}}.$ (4)
2.4 GPI via the Option Keyboard
The Option Keyboard (OK; Barreto et al., , 2019, 2020) is a method that generalizes the GPI policy (Eq. (2)) by allowing the agent to follow a different weight vector at each time step. That is, the OK learns a meta-policy $\omega:\mathcal{S}\to\mathcal{Z}$ that outputs weights, $\bm{\mathbf{z}}\in\mathcal{Z}$, [^1] Without loss of generality, we let $\mathcal{Z}$ be the space of $d$-dimensional unit vectors, i.e., $\mathcal{Z}=\{\bm{\mathbf{z}}\in\mathbb{R}^{d}\mid||\bm{\mathbf{z}}||_{2}=1\}$. that are used when following the GPI policy $\pi^{\text{GPI}}(s,\bm{\mathbf{z}};\Pi)$ at each state $s$. Intuitively, given a set of policies $\Pi$, the policy $\omega$ modulates the behavior induced by $\pi^{\text{GPI}}$ via $\Pi$ by controlling the preferences $\omega(s)$ over features for each state $s$. In this case, the OK policy, $\pi^{\text{OK}}_{\omega}$, at each state $s$, is given by:
$\pi_{\omega}^{\text{OK}}(s;\Pi)\triangleq\underset{a\in\mathcal{A}}{\arg\max}\,\,\underset{\pi\in\Pi}{\max}\,\,\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\omega(s).$ (5)
Also note that $\pi_{\omega}^{\text{OK}}(s;\Pi)=\pi^{\text{GPI}}(s,\omega(s);\Pi)$. Intuitively, in the “keyboard” analogy, $\omega(s)$ (a “chord”) combines base policies $\Pi$ (“keys”) to generate more complex behaviors. A key property of the OK is that learning a policy over $\mathcal{Z}$ is often easier than learning one over the original action space $\mathcal{A}$, as solutions typically allow the same action $\bm{\mathbf{z}}$ to be repeated for multiple timesteps, similar to temporally extended options Sutton et al., (1999).
3 Optimal Behavior Basis
In this section, we formalize the problem of identifying a behavior basis—i.e., a set of base policies—that enables the OK to optimally solve all tasks within a given family of MDPs. We are interested in incremental methods for constructing a behavior basis, $\Pi_{k}$, that provably enables the OK to solve any MDP in $\mathcal{M}$. Specifically, we consider approaches where an agent first iteratively learns a behavior basis for solving some subset of tasks, $\mathcal{M}^{\prime}\subset\mathcal{M}$. Then, the method should be capable of leveraging GPI or OK to identify additional specialized policies for optimally solving novel, unseen tasks $M\not\in\mathcal{M}^{\prime}$. This process should progressively expand the behavior basis until it converges to a small but sufficient set that guarantees zero-shot optimality. The core problem investigated in this paper is, then, how to identify an optimal set of base policies that an agent should learn to facilitate transfer learning within specific classes of MDPs.
Let $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}\subseteq\mathcal{M}$ be the—possibly infinite—set of MDPs associated with all linearly expressible reward functions. This set, typically studied in the SFs literature, can be defined as
$\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}\triangleq\{(\mathcal{S},\mathcal{A},p,r_{\bm{\mathbf{w}}},\mu,\gamma)\mid r_{\bm{\mathbf{w}}}=\bm{\mathbf{\phi}}\cdot\bm{\mathbf{w}}\}.$ (6)
In what follows, we consider weight vectors that induce convex combinations of features; that is, $\mathcal{W}=\{\bm{\mathbf{w}}\mid\sum_{i}w_{i}=1$, $w_{i}\geq 0,\forall i\}$. This is a common practice, e.g., in the multi-objective RL literature (Hayes et al., , 2022), since the optimal policy of any MDP remains unchanged when its reward function is scaled by a positive constant, as in $\mathcal{W}$. Existing algorithms that optimally solve all tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ typically identify a set of policies, $\Pi_{k}=\{\pi_{i}\}_{i=1}^{n}$, such that their associated SF vectors, $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$, form a convex coverage set (CCS; Alegre et al., , 2022). A CCS is a set of SF vectors that allows the optimal value function $v^{*}_{\bm{\mathbf{w}}}=\bm{\mathbf{\psi}}^{\pi^{*}_{\bm{\mathbf{w}}}}\cdot\bm{\mathbf{w}}$ for any given task $\bm{\mathbf{w}}\in\mathcal{W}$ to be directly identified: [^2] A CCS, as defined in Eq. (7), may not be unique since tasks may allow for many optimal policies—each with a possibly different SF vector. In what follows, mentions of a CCS refer specifically to the minimal set satisfying Eq.(7), which is guaranteed to be unique (Roijers et al., , 2013).
$\mathrm{CCS}$$\triangleq\{\bm{\mathbf{\psi}}^{\pi}\ |\ \exists\bm{\mathbf{w}}\in\mathcal{W}\ \text{s.t.}\ \forall\pi^{\prime},\bm{\mathbf{\psi}}^{\pi}\cdot\bm{\mathbf{w}}\geq\bm{\mathbf{\psi}}^{\pi^{\prime}}\cdot\bm{\mathbf{w}}\}.$ (7)
Unfortunately, methods that compute the complete CCS for $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ do not scale to complex tasks and become impractical as the number of reward features $d$ grows. Hence, it becomes critical to develop novel techniques capable of recovering all solutions induced by a CCS without incurring the cost of learning the corresponding full set of policies. This could be achieved, for instance, by methods capable of expressing all solutions in a CCS by combining policies in a behavior basis $\Pi_{k}$ smaller than the CCS; that is, $|\Pi_{k}|\leq|\mathrm{CCS}|$.
We start by extending the definition of the OK (Eq. (5)) to include the description of the task being solved, $\bm{\mathbf{w}}\in\mathcal{W}$, as one of its arguments. That is, we consider meta-policies $\omega:\mathcal{S}\times\mathcal{W}\to\mathcal{Z}$, and OK policies defined as
$\pi^{\text{OK}}_{\omega}(s,\bm{\mathbf{w}};\Pi)\triangleq\operatorname*{arg\,max}_{a\in\mathcal{A}}\max_{\pi\in\Pi}\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\omega(s,\bm{\mathbf{w}}).$ (8)
Intuitively, the meta-policy $\omega(s,\bm{\mathbf{w}})$ enables the OK to decompose the optimal policy for a task $\bm{\mathbf{w}}$ by dynamically assigning state-dependent linear weights to the SFs of its various base policies. Note that if $\omega(s,\bm{\mathbf{w}})=\bm{\mathbf{w}}$ for all $s\in\mathcal{S}$, we recover GPI.
Given an arbitrary set of base policies $\Pi_{k}$, we define the space of all policies expressible by the OK, $\Pi^{\text{OK}}(\Pi_{k})$, and their associated SF vectors, $\Psi^{\text{OK}}(\Pi_{k})$, respectively, as
$\Pi^{\text{OK}}(\Pi_{k})$$\triangleq\{\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi_{k})\mid\bm{\mathbf{w}}\in\mathcal{W},\omega:\mathcal{S}\times\mathcal{W}\to\mathcal{Z}\},$
$\Psi^{\text{OK}}(\Pi_{k})$$\triangleq\{\bm{\mathbf{\psi}}^{\pi}\mid\pi\in\Pi^{\text{OK}}(\Pi_{k})\}.$ (9)
We are now ready to mathematically define our goal:
Goal: Learn a set of policies (the behavior basis) $\Pi_{k}$ and a meta-policy $\omega$ such that (1) $|\Pi_{k}|\leq|\mathrm{CCS}|$; and (2) $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is optimal for any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. The latter condition implies that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$; i.e., the OK is at least as expressive as a CCS.
As observed by Barreto et al., (2019, 2020), learning a meta-policy, $\omega$, is often easier than learning base policies over the original action space, $\mathcal{A}$. Consider, e.g., that if a task $\bm{\mathbf{w}}\in\mathcal{W}$ can be solved by switching between two base policies in $\Pi_{k}$, then an optimal meta-policy for $\bm{\mathbf{w}}$, $\omega(\cdot,\bm{\mathbf{w}})$, only needs to output two vectors $\bm{\mathbf{z}}$ (one for each base policy) for any given state. Thus, solving the goal above is bound to be more efficient than constructing a complete CCS since it requires learning fewer optimal base policies.
4 Constructing an Optimal Behavior Basis
In this section, we introduce Option Keyboard Basis (OKB), a novel method to solve the goal introduced in the previous section. The OKB learns a set of base policies, $\Pi_{k}$, and a meta-policy, $\omega$, such that the induced OK policy $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is provably optimal w.r.t. any given task $\bm{\mathbf{w}}\in\mathcal{W}$. OKB’s pseudocode is shown in Alg. 1.
The algorithm starts with a single base policy—optimized to solve an arbitrary initial task; e.g., $\bm{\mathbf{w}}{=}[1/d,...,1/d]^{\top}$—in its set of policies $\Pi$ (lines 2–3). OKB’s initial partial CCS, $\Psi^{\text{OK}}$, and weight support set, $\mathcal{W}^{\text{sup}}$ are initialized as empty sets (lines 4–5). $\mathcal{W}^{\text{sup}}$ will store the weights of tasks the meta-policy has been trained on. At each iteration $k$, OKB carefully selects the weight vectors on which its meta-policy, $\omega$, is trained so that the policies expressible by $\Psi^{\text{OK}}$ iteratively approximate a CCS. This process is implemented by OK-LS (Alg. 2), an algorithm inspired by the SFs Optimistic Linear Support (SFOLS) method (Alegre et al., , 2022) but that operates over a meta-policy $\omega$ rather than the space of base policies.
Algorithm 1 Option Keyboard Basis (OKB)
1:  Input: MPC $(\mathcal{S},\mathcal{A},p,\gamma,\mu)$ with features $\bm{\mathbf{\phi}}(s,a){\in}\mathbb{R}^{d}$.
[…]
3:  $\Pi_{0}\leftarrow\{\pi_{\bm{\mathbf{w}}}\};\Psi^{\text{base}}_{0}\leftarrow\{\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\}$
4:  $\Psi^{\text{OK}}_{0}\leftarrow\{\};\mathcal{W}^{\text{sup}}_{0}\leftarrow\{\}$
5:  Initialize meta-policy $\omega_{0}$
6:  for $k=0,...,\infty$ do
7:     $\triangleright$ Update meta-policy $\omega$
8:     $\omega_{k+1},\Psi^{\text{OK}}_{k+1},\mathcal{W}^{\text{sup}}_{k+1}\leftarrow\texttt{OK-LS}(\omega_{k},\Psi^{\text{OK}}_{k},\mathcal{W}^{\text{sup}}_{k},\Pi_{k})$
9:     $\mathcal{C}\leftarrow\{\}$
10:     for $\bm{\mathbf{w}}\in\texttt{{CornerW}}(\Psi^{\text{base}}_{k})\cup\texttt{{CornerW}}(\Psi^{\text{OK}}_{k+1})$ do
11:        $\triangleright$ Check if task $\bm{\mathbf{w}}$ is solvable with OK and $\Pi_{k}$
12:        if $\pi^{*}_{\bm{\mathbf{w}}}$ is not in $\Pi^{\textup{OK}}(\Pi_{k})$ then
13:           $\mathcal{C}\leftarrow\mathcal{C}\cup\{\bm{\mathbf{w}}\}$
14:        end if
15:     end for
16:     if $\mathcal{C}$ is empty then
17:        $\triangleright$ Found optimal basis $\Pi$ and meta-policy $\omega$
18:        Return $\omega_{k+1},\Pi_{k},\Psi^{\text{OK}}_{k+1}$
19:     else
20:        $\triangleright$ Learn a new base policy
21:        $\bm{\mathbf{w}}\leftarrow$ select a task from $\mathcal{C}$
22:        $\pi_{\bm{\mathbf{w}}},\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\leftarrow\texttt{NewPolicy}(\bm{\mathbf{w}})$
23:        $\Pi_{k+1},\leftarrow\Pi_{k}\cup\{\pi_{\bm{\mathbf{w}}}\};\Psi^{\text{base}}_{k+1}\leftarrow\Psi^{\text{base}}_{k}\cup\{\bm{\mathbf{\psi}}^{\pi_{\bm{\mathbf{w}}}}\}$
24:        $\Pi_{k+1},\Psi^{\text{base}}_{k+1}{\leftarrow}\texttt{RemoveDominated}(\Pi_{k+1},\Psi^{\text{base}}_{k+1})$
25:     end if
26:  end for
Algorithm 2 OK - Linear Support (OK-LS)
1:  Input: Meta-policy $\omega$, partial CCS $\Psi^{\text{OK}}$, weight support $\mathcal{W}^{\text{sup}}$, base policies $\Pi_{k}$.
2:  while True do
3:     $\mathcal{W}^{\text{corner}}\leftarrow\texttt{CornerW}(\Psi^{\text{OK}})\setminus\mathcal{W}^{\text{sup}}$
4:     if $\mathcal{W}^{\text{cw}}$ is empty then
5:        Return $\omega,\Psi^{\text{OK}},\mathcal{W}^{\text{sup}}$
6:     end if
7:     $\mathcal{W}^{\text{sup}}\leftarrow\mathcal{W}^{\text{sup}}\cup\mathcal{W}^{\text{corner}}$
8:     $\omega\leftarrow\texttt{TrainOK}(\omega,\mathcal{W}^{\text{sup}},\Pi_{k})$triangleright (Alg. 3)
9:     $\Psi^{\text{OK}}\leftarrow\{\bm{\mathbf{\psi}}^{\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi_{k})}\mid\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}\}$
10:     $\Psi^{\text{OK}}\leftarrow\texttt{RemoveDominated}(\Psi^{\text{OK}})$
11:  end while
Next, in lines 10–15, OKB identifies a set of candidate tasks, $\mathcal{C}$, that the OK cannot optimally solve with its current set of policies, $\Pi_{k}$—i.e., tasks for which $\pi_{\bm{\mathbf{w}}}^{*}$ is not included in $\Pi^{\text{OK}}(\Pi_{k})$. We discuss how to check this condition in Section 4.1 If $\mathcal{C}$ is empty (line 16), the OKB terminates and returns a meta-policy ($\omega$) and base policies ($\Pi_{k}$) capable of ensuring that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$. This is due to Thm. 4.3, which we discussed later. If $\mathcal{C}$ is not empty, OKB selects a task from it and adds a new corresponding optimal base policy to its behavior basis, $\Pi_{k}$ (lines 20–23). In line 24, RemoveDominated removes redundant policies—policies that are not strictly required to solve at least one task.
In Algorithms 1 and 2, the function $\texttt{CornerW}(\Psi)$ takes a set of SF vectors, $\Psi$, as input and returns a set of corner weights (see App. E.1 for its mathematical definition). Intuitively, both Alg. 1 and Alg. 2 rely on the fact that (i) a task $\bm{\mathbf{w}}\in\mathcal{W}$ that maximizes $\Delta(\bm{\mathbf{w}})\triangleq v^{*}_{\bm{\mathbf{w}}}-\max_{\pi\in\Pi_{k}}v^{\pi}_{\bm{\mathbf{w}}}$ is guaranteed to be a corner weight, and (ii) if OKB has an optimal policy for all corner weights, then it must have identified a CCS. These results were demonstrated in prior work on constructing a CCS (Roijers et al., , 2013; Alegre et al., , 2022) and can also be proven using the fundamental theorem of linear programming.
OK-LS (Alg. 2) trains the meta-policy $\omega$, on selected tasks $\mathcal{W}^{\text{cw}}$ (line 3) using the base policies $\Pi_{k}$, so that is partial CCS, $\Psi^{\text{OK}}$, iteratively approximates a CCS. It stores the tasks $\bm{\mathbf{w}}$ it has already trained on, as well as the corner weights of the current iteration, in the weight support set, $\mathcal{W}^{\text{sup}}$ (line 7). If $\mathcal{W}^{\text{cw}}$ is empty in a given iteration, no tasks remain to be solved, and the algorithm returns the updated meta-policy. Otherwise, OK-LS adds the corner weights $\mathcal{W}^{\text{cw}}$ to $\mathcal{W}^{\text{sup}}$ and trains the meta-policy $\omega$ on the tasks $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$ using the TrainOK subroutine (Alg. 3). Finally, in line 9, OK-LS computes the SF vectors of the policies induced by OK for each $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$, updating its partial CCS, $\Psi^{\text{OK}}$. [^3] Note that to reduce the algorithm’s computational cost, SF vectors can be trained and computed only for the weight vectors in $\mathcal{W}^{\text{sup}}$ that are new in the current iteration (lines 8–9). In App. C, we discuss how we train $\omega$ using an actor-critic RL method (Alg. 3). Finally, note that to accelerate the learning of $\omega$, the corner weights in $\mathcal{W}^{\text{cw}}$ can be prioritized, similar to Alegre et al., (2022).
4.1 Condition for OK Optimality
In this section, we address the following question: Given an arbitrary task with reward function $r$ and a set of base policies $\Pi_{k}$, is $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$? In other words, does a meta-policy $\omega$ exist such that the OK policy, $\pi^{\text{OK}}{\omega}(\cdot;\Pi_{k})$, can represent the optimal policy $\pi_{r}^{*}$? Determining whether this holds is essential for implementing the OKB step in line 12.
Proposition 4.1.
Let $\Pi_{k}=\{\pi_{i}\}_{i=1}^{n}$ be a set of base policies with corresponding SFs $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$. Given an arbitrary reward function $r$, an optimal OK policy $\pi^{\text{OK}}_{\omega}(\cdot;\Pi)$ can only exist if there exists an OK meta-policy, $\omega:\mathcal{S}\to\mathcal{Z}$, such that for all $s\in\mathcal{S}$,
$\operatorname*{arg\,max}_{a\in\mathcal{A}}\max_{\pi\in\Pi_{k}}\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\omega(s)=\operatorname*{arg\,max}_{a\in\mathcal{A}}q^{*}_{r}(s,a).$ (10)
This proposition provides a sufficient condition for the existence of an optimal OK meta-policy $\omega$ for a given reward function $r$. Intuitively, the OK policy should be able to express all optimal actions through $\omega$ and $\Psi$. Next, we show how to verify this condition without requiring access to $q^{*}_{r}$.
Let $\Pi_{k}$ be a set of base policies, $\omega$ be an OK meta-policy, and $q^{\omega}_{r}(s,\bm{\mathbf{z}})=r(S_{t},A_{t})+\gamma\mathbb{E}_{\omega}[q^{\omega}_{r}(S_{t+1},\omega(S_{t+1}))\mid S_{t}=s,A_{t}=\pi^{\text{GPI}}(S_{t},\bm{\mathbf{z}};\Pi_{k})]$ be the meta-policy’s action-value function for task $r$. Given a state-action pair $(s,a)$, let the advantage function of $\omega$ for executing action $a$ in state $s$ be
$A^{\omega}_{r}(s,a)\triangleq r(s,a)+\gamma\mathbb{E}_{\omega}[q^{\omega}_{r}(S_{t+1},\omega(S_{t+1}))\mid S_{t}{=}s,A_{t}{=}a]-q^{\omega}_{r}(s,\omega(s)).$ (11)
It is well-known that an optimal policy’s advantage function is zero when evaluated at an optimal action; i.e., $A^{\pi^{*}}(s,a^{*})=0$ for all $s\in\mathcal{S}$. Thm. 4.2 uses this insight to introduce a principled way to verify if $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$.
Theorem 4.2.
Let $\Pi_{k}$ be a set of base policies and $\omega^{*}$ be a meta-policy trained to convergence to solve a given task $r$. If $A^{\omega^{*}}{r}(s,a)>0$ for some $(s,a)$, then action $a$ is not expressible by $\pi^{\text{OK}}{\omega^{*}}(\cdot;\Pi_{k})$. Consequently, the OK with base policies $\Pi_{k}$ cannot represent an optimal policy for $r$.
For a detailed discussion on applying Thm. 4.2 in practical implementations of the OKB to verify whether $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$, see App. D.
4.2 Theoretical Results
In this section, we present the theoretical guarantees of OKB (Alg. 1). Proofs of the theorems can be found in App. A.
Theorem 4.3.
OKB returns a set of base policies $\Pi_{k}$ and a meta-policy $\omega$ such that $|\Pi_{k}|\leq|\mathrm{CCS}|$ and $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is optimal for any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. This implies that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$; i.e., the OK is at least as expressive as a CCS.
This theorem states the main result of this paper: the proposed method, OKB (Alg. 1), achieves the primary goal introduced in Section 3. That is, OKB identifies a behavior basis, $\Pi_{k}$, that can be used by an option keyboard to optimally solve any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. Furthermore, since OKB allows optimal policies to be expressed with fewer base policies than a CCS, the total computational cost of identifying optimal solutions for all $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ is significantly reduced compared to GPI-based algorithms. This is because computing all policies needed to form a CCS becomes intractable as the number of reward features $d$ grows. As a result, the computational cost gains provided by OKB become increasingly pronounced as task complexity increases.
To state our next theoretical result, we first recall Thm. 2 by Barreto et al., (2017):
Theorem 4.4 (Barreto et al., (2017)).
Let $\Pi=\{\pi_{i}\}_{i=1}^{n}$ be a set of optimal policies w.r.t. tasks $\{\bm{\mathbf{w}}_{i}\}_{i=1}^{n}$ and let $\{\hat{\bm{\mathbf{\psi}}}^{\pi_{i}}\}_{i=1}^{n}$ be approximations to their SFs. Let $\bm{\mathbf{w}}\in\mathcal{W}$ be a task and let $|q^{\pi_{i}}_{\bm{\mathbf{w}}}(s,a)-\hat{q}^{\pi_{i}}_{\bm{\mathbf{w}}}(s,a)|\leq\epsilon\text{ for all }(s,a)\in\mathcal{S}\times\mathcal{A},\text{ and }\,\pi_{i}\in\Pi$. Let $\bm{\mathbf{\phi}}_{\mathrm{max}}\triangleq\max_{s,a}||\bm{\mathbf{\phi}}(s,a)||$. Then, it holds that, for all $(s,a)\in\mathcal{S}\times\mathcal{A}$:
$q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{GPI}}}_{\bm{\mathbf{w}}}(s,a)\leq\frac{2}{1-\gamma}\left(\bm{\mathbf{\phi}}_{\mathrm{max}}\min_{i}||\bm{\mathbf{w}}-\bm{\mathbf{w}}_{i}||+\epsilon\right).$
This theorem describes the optimality gap of GPI policies, how it depends on the available base policies, and how it is affected by function approximation errors. It does not, however, characterize the optimality gap of option keyboards, which generalize GPI policies. We introduce and highlight two generalizations of this theorem: (1) Since the OK generalizes GPI policies, it follows that $q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{OK}}_{\omega}}_{\bm{\mathbf{w}}}(s,a)\leq q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{GPI}}}_{\bm{\mathbf{w}}}(s,a)$; and (2) when Eq. (10) (Prop. 4.1) is satisfied, $q^{*}_{\bm{\mathbf{w}}}(s,a)-q^{\pi^{\text{OK}}_{\omega}}_{\bm{\mathbf{w}}}(s,a)=0$. That is, the OK can completely avoid optimality gaps due to approximation errors in the base policies’ SFs. Hence, not only are the OK and OKB more expressive than GPI (Thm. 4.3), but they also naturally lead to transfer learning strategies that are significantly more robust to approximation errors than GPI.
Theorem 4.5.
Let $\Pi_{k}$ be the set of policies learned by OKB (Alg. 1) when solving tasks that are linearly expressible in terms of the reward features $\bm{\mathbf{\phi}}(s,a)\in\mathbb{R}^{d}$. Let $r$ be an arbitrary reward function that is non-linear with respect to $\bm{\mathbf{\phi}}$. Suppose the optimal policy for $r$ can be expressed by alternating, as a function of the state, between policies in the CCS induced by $\bm{\mathbf{\phi}}$. That is, suppose that for all $s\in\mathcal{S}$, there exists a policy $\pi^{*}_{\bm{\mathbf{w}}}$ (optimal for some $\bm{\mathbf{w}}\in\mathcal{W}$) such that $\pi^{*}_{r}(s)=\pi^{*}_{\bm{\mathbf{w}}}(s)$. Then, the OK can represent the optimal policy for $r$ using the set of base policies $\Pi_{k}$; i.e., $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$.
This theorem precisely characterizes a class of non-linear tasks that can be optimally solved by an OK using the behavior basis learned by OKB. Intuitively, OKB can solve any task whose optimal policy can be constructed by alternating between the optimal policies for tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$, even when the task itself is non-linear. Importantly, existing methods that compute a CCS can only handle linearly expressible tasks, whereas OKB extends this capability to a broader class of problems. Furthermore, OKB achieves this without having to learn all policies in the CCS (Thm. 4.3). We further discuss the properties of OK under non-linear reward functions in App. B.
5 Experiments
In this section, we empirically evaluate OKB and investigate the following research questions: Q1: Can OKB approximate a CCS more effectively while requiring fewer base policies than competing methods? Q2: Does OKB’s performance advantage become more pronounced as problem complexity increases, i.e., as the number of reward features $d$ grows? Q3: Can the base policies learned by OKB be leveraged to solve tasks with non-linear reward functions under the conditions outlined in Thm. 4.5?
Figure 1: Domains used in the experiments: Minecart, FetchPickAndPlace, Item Collection, and Highway.
Fig. 1 depicts the domains used in our experiments. To handle the high-dimensional state space of these domains, we learn base policies using a Universal SF Approximator (USFA) Borsa et al., (2019). Furthermore, rather than learning a separate SF for each base policy $\pi\in\Pi_{k}$, we train a single USFA, $\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})$, conditioned on task vectors $\bm{\mathbf{w}}$, such that $\pi_{\bm{\mathbf{w}}}(s)\approx\operatorname*{arg\,max}_{a\in\mathcal{A}}\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})\cdot\bm{\mathbf{w}}$. At each call to NewPolicy($\bm{\mathbf{w}}$) in Alg.1, the USFA is trained to solve the new task $\bm{\mathbf{w}}$. Additional experimental details are provided in App. E.
(a) Minecart domain.
(b) Highway domain.
Figure 2: Mean normalized return per iteration for each method on a set of test tasks in the (a) Minecart and (b) Highway domains.
Figure 3: Mean normalized return over test tasks as a function of iteration number—i.e., number of base policies learned per method [FetchPickAndPlace]. As $d$ increases, OKB’s performance advantage over SFOLS (a state-of-the-art GPI-based algorithm) grows.
In all experiments, we report the mean normalized return of each method (normalized with respect to the minimum and maximum returns observed for a given task) along with the 95% bootstrapped confidence interval over 15 random seeds. We compare OKB to SFOLS (Alegre et al., , 2022), a method that identifies a CCS using GPI policies, and to OKB-Uniform, a variant of OKB that selects tasks uniformly from $\mathcal{W}$ in line 21 of Alg. 1, rather than from $\mathcal{C}$. At each iteration, all methods have a fixed budget of environment interactions. To ensure fair comparisons, since OKB must also train a meta-policy while SFOLS does not, we restrict OKB to using only half of its budget for learning base policies, allocating the other half to training the meta-policy (Alg. 2). This makes the comparison more conservative for OKB, as it has fewer interactions available to learn base policies.
In Fig.2, we show the mean return of each method over a set of test tasks, $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$, [^4] To generate test task sets $\mathcal{W}^{\prime}\subset\mathcal{W}$ for different values of $d$, we used the method introduced by Takagi et al., (2020), which produces uniformly spaced weight vectors in $\mathcal{W}$. as a function of the iteration number (i.e., the number of base policies learned). We report results for the Minecart domain—a classic multi-objective RL problem—and the Highway domain. Both domains have reward functions defined by $d=3$ reward features, which are detailed in App. E.2. These results demonstrate that OKB achieves strong performance across test tasks with a small number of base policies, positively answering research question Q1: OKB can approximate a CCS more effectively using fewer base policies. This is particularly evident in Fig. 2(a), where OKB reaches near-optimal performance with just 2–3 base policies. While OKB-Uniform also performs well in the Minecart domain, its lower performance in the Highway domain (Fig. 2(b)) highlights the importance of expanding the set of base policies by carefully selecting promising tasks—defined by corner weights—for training. Across both domains, OKB consistently outperforms all competing methods.
To investigate Q2, we evaluate each method in the FetchPickAndPlace domain with varying numbers of reward features, $d\in\{2,4,6,8\}$. In this domain, an agent controls a robotic arm that must grasp a block from a table and move it to a specified location. Each of the $d$ reward features represents the block’s distance to a different target location (shown in red in Fig. 1). Fig. 3 shows that as the number of target locations ($d$) increases, the performance gap between OKB and SFOLS increases significantly—positively answering Q2. Intuitively, OKB focuses on tasks where the OK cannot express optimal actions (see Thm. 4.2). By learning the corresponding optimal policies, OKB quickly identifies a behavior basis that enables the OK to solve tasks across the entire space of task vectors $\mathcal{W}$. Conversely, the gap between OKB and OKB-Uniform decreases since a larger number of reward features enhances the expressivity of the OK (Prop. 4.1). The total computation time required by OKB to train a meta-policy and base policies to ensure zero-shot optimality is consistently comparable to or lower than that of competing methods. This is because OKB requires fewer policies to be learned and evaluated at inference time.
Figure 4: Mean total reward in the Item Collection domain under a non-linear reward function.
Figure 5: Continuous actions produced by the meta-policy while solving a sample task. The color of each timestep’s column indicates the selected base policy. Notice OKB selects base policies in a temporally consistent manner, suggesting these policies encode temporally-extended behaviors useful to solving the task.
Next, we investigate Q3 by conducting an experiment similar to the one proposed by Alver and Precup, (2022). We compare OKB to relevant competitors in an environment with non-linear reward functions. The Item Collection domain (Fig. 1; top right) is a $10\times 10$ grid world where agents must collect two types of items. Reward features are indicator functions that signal whether the agent has collected a particular item in the current state. This domain requires function approximation due to the combinatorial number of possible states. After OKB learns a behavior basis $\Pi_{k}$, the agent trains a meta-policy $\omega:\mathcal{S}\to\mathcal{Z}$ to solve a task with a non-linear reward function. Specifically, this is a sequential task where the agent must collect all instances of one item type before collecting any items of another type. We compare OKB with SIP (Alver and Precup, , 2022), [^5] SIP was not included in previous experiments since it assumes independent reward features (Alver and Precup, , 2022). a method that learns $d$ base policies—each maximizing a specific reward feature—and a meta-policy $\omega$ over a discrete set of weights in $\mathcal{W}$. We also compare OKB against a baseline DQN agent that learns tasks from scratch and against GPI policies (horizontal blue lines), which are optimal for either exclusively prioritizing one item type or assigning equal importance to both. Fig. 4 shows the mean total reward achieved by each method in this non-linear task as a function of the total number of environment interactions required to train the meta-policies. Both OKB and SIP successfully solved the task, with OKB doing so more quickly. The DQN baseline failed to solve the task as directly learning a policy over $\mathcal{A}$ is significantly more difficult.
Finally, we examine the qualitative behavior of OKB policies by visualizing the continuous actions produced by the meta-policy $\omega(s,\bm{\mathbf{w}})$ while solving a sample task in the Highway domain, where the agent must prioritize driving fast and staying in the rightmost lane. In Fig. 5, the color of each timestep’s column represents the selected base policy. The agent initially accelerates as quickly as possible by following base policy $\pi_{1}$ for the first 10 timesteps while the rightmost lane is occupied. At timestep $t=10$, the agent turns right and switches to base policy $\pi_{2}$, which is optimized for driving in the rightmost lane. At $t=20$, $\pi_{1}$ becomes active again, allowing the agent to accelerate while staying in the rightmost lane. Notice that while the meta-policy $\omega$ is relatively smooth, the base policies it selects are complex. In particular, although $\bm{\mathbf{z}}_{t}$ remains approximately constant during the first 10 timesteps, the underlying primitive actions $a\in\mathcal{A}$ (Fig. 5; black numbers at the top of each column) continuously alternate between turning left, turning right, accelerating, and idling. Finally, notice that OKB selects base policies in a temporally consistent manner, suggesting that it learns to identify temporally extended behaviors—akin to options Sutton et al., (1999)—that help solve the task.
6 Related Work
OK and GPI.
Previous works have extended the OK in different ways. Carvalho et al., 2023b introduced a neural network architecture for jointly learning reward features and SFs, while Machado et al., (2023) proposed using the OK with base policies that maximize Laplacian-based reward features. However, unlike our work, these methods do not provide theoretical guarantees on optimality for solving specific families of tasks. Other works have introduced GPI variants for risk-aware transfer (Gimelfarb et al., , 2021), mitigating function approximation errors (Kim et al., , 2024), planning with approximate models (Alegre et al., 2023a, ), and combining non-Markovian policies (Thakoor et al., , 2022). These approaches improve GPI in ways that are orthogonal to our contribution and could potentially be combined with our method.
Learning behavior basis.
Previous works have addressed the problem of learning an effective behavior basis for transfer learning within the SF framework (Nemecek and Parr, , 2021; Zahavy et al., , 2021). The method introduced by Alver and Precup, (2022) assumes that reward features are independent (i.e., can be controlled independently) and that the MDP’s transition function is deterministic. These assumptions often do not hold in complex RL domains, including most of those studied in Section 5. Alegre et al., (2022) proposed a method that learns a set of policies corresponding to a CCS and combines them with GPI. In contrast, we use option keyboards, which enable a broader range of policies to be expressed, allowing our method to approximate a CCS more effectively with a smaller behavior basis.
Learning features.
While we focused on identifying an optimal behavior basis with respect to a given set of reward features, other works have addressed the complementary problem of learning more expressive reward features (Touati and Ollivier, , 2021; Carvalho et al., 2023a, ; Chua et al., , 2024). A promising future research direction is to combine OKB with methods such as the Forward-Backward representation Touati et al., (2023) to construct an optimal behavior basis under learned reward features.
7 Conclusions
We introduced OKB, a principled method with strong theoretical guarantees for identifying the optimal behavior basis for the Option Keyboard (OK). Our theoretical results, supported by thorough empirical analysis, show that OKB significantly reduces the number of base policies required to achieve zero-shot optimality in new tasks compared to state-of-the-art methods. In particular, OKB constructs a behavior basis efficiently by carefully selecting base policies that iteratively improve its meta-policy’s approximation of the CCS. We empirically evaluate OKB in challenging high-dimensional RL problems, demonstrating that it consistently outperforms GPI-based approaches. Notably, its performance advantage becomes increasingly pronounced as the number of reward features grows. Finally, we prove that OKB’s expressivity surpasses that of a CCS, enabling it to optimally solve specific classes of non-linear tasks.
An interesting direction for future research is extending OKB with temporally extended meta-policies $\omega$ that incorporate learned termination conditions. Additionally, given the recently identified mathematical connections between multi-objective RL and SFs (Alegre et al., , 2022), OKB could be adapted and combined with existing MORL techniques to further improve sample complexity.
Appendix A Proofs
A.1 Proof of Proposition 4.1
Proposition (4.1).
Let $\Pi_{k}=\{\pi_{i}\}_{i=1}^{n}$ be a set of base policies with corresponding SFs $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$. Given an arbitrary reward function $r$, an optimal OK policy $\pi^{\text{OK}}_{\omega}(\cdot;\Pi)$ can only exist if there exists an OK meta-policy, $\omega:\mathcal{S}\to\mathcal{Z}$, such that for all $s\in\mathcal{S}$,
$\operatorname*{arg\,max}_{a\in\mathcal{A}}\max_{\pi\in\Pi_{k}}\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\omega(s)=\operatorname*{arg\,max}_{a\in\mathcal{A}}q^{*}_{r}(s,a).$ (12)
Proof.
The proof follows directly from the definition of a deterministic optimal policy for a given reward function $r$. Recall that $\pi^{*}_{r}(s)\in\operatorname*{arg\,max}_{a\in\mathcal{A}}q^{*}_{r}(s,a)$ and $\pi^{\text{OK}}_{\omega}(s;\Pi)\in\operatorname*{arg\,max}_{a\in\mathcal{A}}\max_{\pi\in\Pi}\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\omega(s)$. If for some state $s\in\mathcal{S}$, there is no value of $\bm{\mathbf{z}}=\omega(s)$ such that Eq. (12) holds, then it is not possible to define a meta policy $\omega$ such that $\pi^{\text{OK}}_{\omega}(s;\Pi)=\pi^{*}_{r}(s)$. ∎
A.2 Proof of Theorem. 4.2
Theorem (4.2).
Let $\Pi_{k}$ be a set of base policies and $\omega^{*}$ be a meta-policy trained to convergence to solve a given task $r$. If $A^{\omega^{*}}{r}(s,a)>0$ for some $(s,a)$, then action $a$ is not expressible by $\pi^{\text{OK}}{\omega^{*}}(\cdot;\Pi_{k})$. Consequently, the OK with base policies $\Pi_{k}$ cannot represent an optimal policy for $r$.
Proof.
First, recall the definitions of $q^{\omega}_{r}(s,\bm{\mathbf{z}})$ and $A^{\omega}_{r}(s,a)$:
$q^{\omega}_{r}(s,\bm{\mathbf{z}})$ $\triangleq r(S_{t},A_{t})+\gamma\mathbb{E}_{\omega}[q^{\omega}_{r}(S_{t+1},\omega(S_{t+1}))\mid S_{t}=s,A_{t}=\pi^{\text{GPI}}(S_{t},\bm{\mathbf{z}};\Pi_{k})],$ (13)
$A^{\omega}_{r}(s,a)$ $\triangleq r(S_{t},A_{t})+\gamma\mathbb{E}_{\omega}[q^{\omega}_{r}(S_{t+1},\omega(S_{t+1}))\mid S_{t}=s,A_{t}=a]-q^{\omega}_{r}(s,\omega(s)).$ (14)
$A^{\omega}_{r}(s,a)$ measures the relative benefit of executing action $a$ in state $s$ compared to following the current OK policy $\pi^{\text{OK}}_{\omega}$.
If $A^{\omega}_{r}(s,a)>0$, it means executing $a$ in $s$ provides a higher expected return than the action currently selected by $\pi^{\text{OK}}_{\omega}$. Hence, if there exists an $(s,a)$ such that $A^{\omega}_{r}(s,a)>0$, it implies that the OK policy fails to represent at least one optimal action, meaning it is suboptimal.
Recall from Prop. 4.1 that for the OK policy to be optimal, it must express the optimal action $a^{*}$ for every $s$:
$\arg\max_{a\in A}\max_{\pi\in\Pi_{k}}\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\omega(s)=\arg\max_{a\in A}q^{*}_{r}(s,a).$
Since $\omega^{*}$ is trained until convergence to solve $r$, it has reached a stable policy. However, if there exists $(s,a)$ such that $A^{\omega^{*}}_{r}(s,a)>0$, then:
$q^{*}_{r}(s,a)>q^{\omega^{*}}_{r}(s,\omega^{*}(s)).$
This means that in at least one state-action pair, the OK policy does not select an action that maximizes the expected return.
Since $\pi^{\text{OK}}_{\omega^{*}}$ is restricted to the actions expressible via its base policies $\Pi_{k}$, this implies that:
$a^{*}\not\in\left\{\arg\max_{a}\max_{\pi\in\Pi_{k}}\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\omega^{*}(s)\right\}.$
Thus, no possible weighting of the base policies can recover the optimal action in state $s$.
Since at least one optimal action $a^{*}$ is missing from the action space of the OK policy, it follows that: $\pi^{OK}_{\omega^{*}}$ is not an optimal policy for $r$, and the task $r$ cannot be optimally solved using the given base policies $\Pi_{k}$. Therefore, to guarantee optimality, additional base policies must be introduced into $\Pi_{k}$. ∎
A.3 Proof of Theorem. 4.3
Lemma A.1 (Alegre et al., (2022)).
Let $\Psi_{k}\subseteq\mathrm{CCS}$ be a subset of the CCS. If, for all corner weights $\bm{\mathbf{w}}\in\texttt{{CornerW}}(\Psi_{k})$, it holds that $\max_{\bm{\mathbf{\psi}}^{\pi}\in\Psi_{k}}\bm{\mathbf{\psi}}^{\pi}\cdot\bm{\mathbf{w}}=v^{*}_{\bm{\mathbf{w}}}$, then $\Psi_{k}$ contains all elements of the CCS, i.e., $\mathrm{CCS}\subseteq\Psi_{k}$.
The result in this lemma follows directly from the theoretical guarantees of previous methods that construct a CCS, e.g., OLS and SFOLS (Roijers, , 2016; Alegre et al., , 2022). Intuitively, this result states that to check if a given SF vector set $\Psi_{k}$ is a CCS, it is enough to check if it contains an optimal policy for all of its corner weights. If this is the case, then it is not possible to identify some alternative weight vector $\bm{\mathbf{w}}\in\mathcal{W}$ for which it does not contain an optimal policy.
Theorem (4.3).
OKB returns a set of base policies $\Pi_{k}$ and a meta-policy $\omega$ such that $|\Pi_{k}|\leq|\mathrm{CCS}|$ and $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is optimal for any task $M\in\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$. This implies that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}(\Pi_{k})$; i.e., the OK is at least as expressive as a CCS.
Proof.
Assumption 1. NewPolicy($\bm{\mathbf{w}}$) returns an optimal policy $\pi_{\bm{\mathbf{w}}}^{*}$ for the given task $\bm{\mathbf{w}}$.
Assumption 2. For all $\bm{\mathbf{w}}\in\mathcal{W}^{\text{sup}}$, if an optimal policy $\pi_{\bm{\mathbf{w}}}^{*}\in\Pi^{\text{OK}}(\Pi)$, then TrainOK($\omega,\mathcal{W}^{\text{sup}},\Pi$) (Alg. 3) returns a meta-policy $\omega$ s.t. $\pi^{\text{OK}}_{\omega}(\cdot,\bm{\mathbf{w}};\Pi)$ is optimal w.r.t. task $\bm{\mathbf{w}}$.
OKB (Alg. 1) maintains two partial CCSs, $\Psi^{\text{base}}_{k}$ and $\Psi^{\text{OK}}_{k}$, respectively being the SFs of the base policies $\Pi_{k}$ and the policies generated via the OK using $\Pi_{k}$. We first will prove that, if in any given iteration $k$, the set of candidate corner weights $\mathcal{C}$ is empty (see line 16 in Alg. 1), then it holds that $\mathrm{CCS}\subseteq\Psi^{\text{OK}}_{k+1}$. First, note that $\mathcal{C}$ contains corner weights of both $\Psi^{\text{base}}_{k}$ and $\Psi^{\text{OK}}_{k}$. If the OK policy $\pi^{\text{OK}}_{\omega}(\cdot;\Pi_{k})$ is able to optimally solve all tasks in $\mathcal{C}$ (line 12 in Alg. 1), then we have that $\mathrm{CCS}\in\Psi^{\text{OK}}(\Pi_{k})$ due to Lemma A.1. We note that assumptions 1 and 2 above are necessary because Lemma A.1 requires the partial CCS $\Psi_{k}$ to be a subset of the CCS. If it contains $\epsilon$-optimal policies, then it is possible to prove convergence to $\epsilon$-CCSs instead (Alegre et al., 2023b, ).
Now, to conclude the proof, we only need to show that $|\Pi_{k}|\leq|\mathrm{CCS}|$. The set of base policies is never larger than the CCS due to line 24 in Algorithm 1, in which dominated base policies (i.e., redundant policies that are not exclusively optimal w.r.t. any $\bm{\mathbf{w}}$) are removed from $\Pi_{k}$ by the procedure RemoveDominated. Importantly, the set of base policies returned by OKB, $\Pi_{k}$, will only have the same of the CCS ($|\Pi_{k}|=|\mathrm{CCS}|$) in domains where no single task in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$ can be optimally solved by combining policies optimal for other tasks in $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}$.  [^6] An example of such a domain is an MDP composed of independent corridors, in which optimal policies for different weights $\bm{\mathbf{w}}\in\mathcal{W}$ must traverse different corridors.
A.4 Proof of Theorem. 4.5
Theorem (4.5).
Let $\Pi_{k}$ be the set of policies learned by OKB (Alg. 1) when solving tasks that are linearly expressible in terms of the reward features $\bm{\mathbf{\phi}}(s,a)\in\mathbb{R}^{d}$. Let $r$ be an arbitrary reward function that is non-linear with respect to $\bm{\mathbf{\phi}}$. Suppose the optimal policy for $r$ can be expressed by alternating, as a function of the state, between policies in the CCS induced by $\bm{\mathbf{\phi}}$. That is, suppose that for all $s\in\mathcal{S}$, there exists a policy $\pi^{*}_{\bm{\mathbf{w}}}$ (optimal for some $\bm{\mathbf{w}}\in\mathcal{W}$) such that $\pi^{*}_{r}(s)=\pi^{*}_{\bm{\mathbf{w}}}(s)$. Then, the OK can represent the optimal policy for $r$ using the set of base policies $\Pi_{k}$; i.e., $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$.
Proof.
We have that, for all $s\in\mathcal{S}$, $\exists\bm{\mathbf{w}}\in\mathcal{W}$ such that:
$\pi^{*}_{r}(s)$ $=\pi^{*}_{\bm{\mathbf{w}}}(s),\quad\text{where }\bm{\mathbf{\psi}}^{\pi^{*}_{\bm{\mathbf{w}}}}\in\mathrm{CCS},$ (15)
$=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bm{\mathbf{\psi}}^{\pi^{*}_{\bm{\mathbf{w}}}}(s,a)\cdot\bm{\mathbf{w}}.$ (16)
Due to Thm. 4.3, we know that by employing the set of base policies returned by OKB, we have that $\pi^{*}_{\bm{\mathbf{w}}}\in\Pi^{\textup{OK}}(\Pi_{k})$. Equivalently,
$\pi^{*}_{\bm{\mathbf{w}}}(s)=\pi^{\text{OK}}_{\omega}(s,\bm{\mathbf{w}};\Pi_{k})=\operatorname*{arg\,max}_{a\in\mathcal{A}}\max_{\pi\in\Pi_{k}}\bm{\mathbf{\psi}}^{\pi}(s,a)\cdot\omega(s,\bm{\mathbf{w}}).$ (17)
Combining Eq. (16) and Eq. (17), we have that
$\pi^{*}_{r}(s)=\pi^{\text{OK}}_{\omega}(s,\bm{\mathbf{w}};\Pi_{k}),$ (18)
which concludes the proof, demonstrating that $\pi^{*}_{r}\in\Pi^{\textup{OK}}(\Pi_{k})$. ∎
Appendix B Beyond Linear Rewards
In this section, we further discuss the use of the OK to solve tasks defined by non-linear reward functions; that is, reward functions that are not linear-expressible under reward features $\bm{\mathbf{\phi}}(s,a,s^{\prime})\in\mathbb{R}^{d}$.
We start by arguing that, in many cases, the linear reward assumption does not pose any limitations. As discussed by Barreto et al., (2020), when both $\mathcal{S}$ and $\mathcal{A}$ are finite, we can recover any possible reward function by defining $d=|\mathcal{S}|^{2}\times|\mathcal{A}|$ features $\phi_{i}$, each being an indicator function associated with a specific transition $(s,a,s^{\prime})$. This result shows that it is possible to define reward features $\bm{\mathbf{\phi}}\in\mathbb{R}^{|\mathcal{S}|^{2}\times|\mathcal{A}|}$, such that every task in $\mathcal{M}$ is linearly expressible. A challenge, however, is to find alternative features with this property, but with dimension $d\ll|\mathcal{S}|$ Touati and Ollivier, (2021).
When examining how the OK is defined (Eq. 5), a natural idea is to use it to solve tasks defined by reward functions with state-dependent weights $\bm{\mathbf{w}}(s)$. In particular, let $\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ be a family of non-linear tasks defined as
$\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}\triangleq\{(\mathcal{S},\mathcal{A},p,r_{\bm{\mathbf{w}}},\mu,\gamma)\mid r_{\bm{\mathbf{w}}}(s,a,s^{\prime})=\bm{\mathbf{\phi}}(s,a,s^{\prime})\cdot\bm{\mathbf{w}}(s)\}.$ (19)
Note that it is easy to show that $\mathcal{M}_{\text{lin}}^{\bm{\mathbf{\phi}}}\subseteq\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}\subseteq\mathcal{M}$.
Unfortunately, the OK may be unable to solve all tasks in the family $\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$, even when having access to a set of policies $\Pi_{k}$ forming a CCS.
Proposition B.1.
There exists a task $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ such that no meta-policy $\omega:\mathcal{S}\to\mathcal{Z}$ and set of policies $\Pi$ results in $\pi^{\text{OK}}_{\omega}(s;\Pi)$ being optimal with respect to $M$.
Proof.
Assume a family of MDPs $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ with states, actions, transitions, and features defined as in Figure 6.
Figure 6: Counterexample MDP.
Let a task $M\in\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ be defined via the following reward function: $r(s_{1})=\bm{\mathbf{\phi}}(s_{1})\cdot[1,0]=0,r(s_{2})=\bm{\mathbf{\phi}}(s_{2})\cdot[-1,-1]=-3,r(s_{3})=\bm{\mathbf{\phi}}(s_{3})\cdot[-1,-1]=-3,r(s_{4})=\bm{\mathbf{\phi}}(s_{4})\cdot[1,1]=2$, with optimal policy $\pi^{*}(s_{0})=a_{4}$ whose SF vector is $\bm{\mathbf{\psi}}^{\pi^{*}}=[1,1]$. We now show that there is no meta-policy $\omega(s)$ that results in $\pi^{\text{GPI}}(s;\omega(s))$ being equal to the optimal policy $\pi^{*}$. This is proved by noticing the impossibility of the following equality:
$\pi^{\text{GPI}}(s_{0},\omega(s_{0}))=\underset{a\in\{a_{1},a_{2},a_{3},a_{4}\}}{\arg\max}\,\,\underset{\pi\in\Pi}{\max}\,\,\bm{\mathbf{\psi}}^{\pi}(s_{0},a)\cdot\omega(s_{0})=a_{4}.$ (20)
Notice that the result above implies that there are tasks in $\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ that can not be solved by $\pi^{\text{OK}}$ even when assuming access to a CCS.
While this result may be negative, we highlight that $\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ is a family of tasks more general than it may seem at first glance.
Proposition B.2.
Let $\mathcal{M}^{\prime}\subset\mathcal{M}$ such that for all $M\in\mathcal{M}^{\prime}$, $r(s,a,s^{\prime})=r(s)$. If $\bm{\mathbf{\phi}}(s)\neq\bm{\mathbf{0}}$ for all $s\in\mathcal{S}$, then $\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}=\mathcal{M}^{\prime}$.
Proof.
If $\bm{\mathbf{\phi}}(s)\neq\bm{\mathbf{0}}$ for all $s\in\mathcal{S}$, then for any reward function $r$ there exist a function $\bm{\mathbf{w}}(s)$ such that $r(s)=\bm{\mathbf{\phi}}(s)\cdot\bm{\mathbf{w}}(s)$. ∎
The proposition above implies that, under mild conditions, $\mathcal{M}_{\text{sd-linear}}^{\bm{\mathbf{\phi}}}$ is able to represent all reward functions and tasks in $\mathcal{M}$.
Appendix C Training the OK Meta-Policy
In Alg. 3, we show the pseudocode for the algorithm that optimizes the OKB meta-policy $\omega$ for a given set of tasks $\mathcal{W}^{\text{sup}}$.
Algorithm 3 Train Option Keyboard (TrainOK)
1:  Input: Meta-policy $\omega$, weight support $\mathcal{W}^{\text{sup}}$, base policies $\Pi$.
2:  Initialize replay buffer $\mathcal{B}$
3:  Let $\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})$ be the critic of the meta-policy $\omega$
4:  $\bm{\mathbf{w}}_{0}\sim\mathcal{W}^{\text{sup}}$; $S_{0}\sim\mu$
5:  for $t=0,...,T$ do
6:     if $S_{t}$ is terminal then
7:        $\bm{\mathbf{w}}_{t}\sim\mathcal{W}^{\text{sup}}$
8:        $S_{t}\sim\mu$
9:     end if
10:     $\bm{\mathbf{z}}_{t}\leftarrow\omega(S_{t},\bm{\mathbf{w}})$
11:     $\bm{\mathbf{z}}_{t}\leftarrow\bm{\mathbf{z}}_{t}+\mathrm{clip}(\epsilon,-0.5,0.5)$, where $\epsilon\sim\mathcal{N}(0,0.2^{2})$  $\triangleright$ Exploration clipped Gaussian noise, as in TD3
12:     $\bm{\mathbf{z}}_{t}\leftarrow\bm{\mathbf{z}}_{t}/||\bm{\mathbf{z}}_{t}||_{2}$
13:     $A_{t}\leftarrow\pi^{\text{GPI}}(S_{t},\bm{\mathbf{z}}_{t};\Pi)$  $\triangleright$ Follow OK policy
14:     Execute $A_{t}$, observe $S_{t+1}$, and $\bm{\mathbf{\phi}}_{t}$
15:     Add $(S_{t},\bm{\mathbf{z}}_{t},\bm{\mathbf{\phi}}_{t},S_{t+1})$ to $\mathcal{B}$
16:     Update $\bm{\mathbf{\psi}}^{\omega}$ by minimizing $\mathbb{E}_{(s,\bm{\mathbf{z}},\bm{\mathbf{\phi}},s^{\prime})\sim\mathcal{B},\bm{\mathbf{w}}\sim\mathcal{W}^{\text{sup}}}\left[\left(\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})-(\bm{\mathbf{\phi}}+\gamma\bm{\mathbf{\psi}}^{\omega}(s^{\prime},\omega(s^{\prime},\bm{\mathbf{w}}),\bm{\mathbf{w}}))\right)^{2}\right]$
17:     Update $\omega$ via the policy gradient $\nabla_{\bm{\mathbf{z}}}\bm{\mathbf{\psi}}^{\omega}(s,\bm{\mathbf{z}},\bm{\mathbf{w}})\cdot\bm{\mathbf{w}}|_{\bm{\mathbf{z}}=\omega(s,\bm{\mathbf{w}})}\nabla_{\omega}\omega(s,\bm{\mathbf{w}})$
18:  end for
19:  Return $\omega$
Appendix D Checking the Condition for OK Optimality
Our practical implementation of the OKB test the condition in Thm. 4.2 by randomly sampling from a replay buffer, =$\mathcal{B}{=}\{(s_{i},a_{i},\bm{\mathbf{\phi}}(s_{i},a_{i}),s^{\prime}_{i},)\}_{i=1}^{n}$, which contains experiences observed during the training of the base policies in $\Pi_{k}$. Given a candidate task $\bm{\mathbf{w}}\in\mathcal{C}$ (line 12 of Alg. 1), we compute the mean positive advantage of $\omega$ as:
$\left(\sum_{i=1}^{|\mathcal{B}|}\mathbbm{1}_{\{\hat{A}_{\bm{\mathbf{w}}}^{\omega}(s_{i},a_{i})>0\}}\right)^{-1}\sum_{i=1}^{|\mathcal{B}|}\max(\hat{A}_{\bm{\mathbf{w}}}^{\omega}(s_{i},a_{i}),0),$ (21)
where $\hat{A}_{\bm{\mathbf{w}}}^{\omega}(s_{i},a_{i})=\bm{\mathbf{\phi}}(s_{i},a_{i})\cdot\bm{\mathbf{w}}+\gamma\bm{\mathbf{\psi}}^{\omega}(s^{\prime}_{i},\omega(s^{\prime}_{i}))\cdot\bm{\mathbf{w}}-\bm{\mathbf{\psi}}^{\omega}(s_{i},\omega(s_{i}))\cdot\bm{\mathbf{w}}$. In line 21 of Alg. 1, we select the task $\bm{\mathbf{w}}\in\mathcal{C}$ with the highest value for Eq. (21). Intuitively, we select the task $\bm{\mathbf{w}}$ for which the OK can improve the most its performance. By adding $\pi_{\bm{\mathbf{w}}}$ to $\Pi_{k}$ and retraining $\omega$, the OK will be more expressive in the subsequent iteration. We highlight that OKB theoretical guarantees (e.g., Thm. 4.3) are independent of the heuristic used to select the task $\bm{\mathbf{w}}\in\mathcal{C}$ (line 21), and other strategies could be used instead.
Appendix E Experiments Details
The code and scripts necessary to reproduce our experiments will be made publicly available upon acceptance.
The USFA $\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})$ used for encoding the base policies $\Pi_{k}$ were modeled with multi-layer perceptron (MLP) neural networks with 4 layers with 256 neurons. We use an ensemble of $10$ neural networks, similar to Chen et al., (2021), and compute the minimum value over two randomly sampled elements when computing the Bellman update targets. We used Leaky ReLU activation functions and layer normalization for improved training stability. We used the same neural network architectures and hyperparameters for OKB and the competing baselines for fairness of comparison.
The budget of environment interactions per iteration (i.e., call to NewPolicy in Alg. 1) used was $25000$, $50000$, $50000$ and $100000$ for the Minecart, FetchPickAndPlace, Item Collection, and Highway domains, respectively. At each iteration, $\bm{\mathbf{\psi}}(s,a,\bm{\mathbf{w}})$ is trained with the current task $\bm{\mathbf{w}}$, as well as with the tasks from previous iterations in order to avoid catastrophic forgetting.
The meta-policy $\omega(s,\bm{\mathbf{w}})$ was modeled with an MLP with 3 layers with 256 neurons. We employed the techniques introduced by Bhatt et al., (2024), i.e., batch normalization and removal of target networks, which increased the training efficiency. We used Adam (Kingma and Ba, , 2015) as the first-order optimizer used to train all neural networks with mini-batches of size $256$.
When training the base policies, we used $\varepsilon$-greedy exploration with a linearly decaying schedule. For training the meta-policy, we added a clipped Gaussian noise (see line 11 of Alg. 3), as done in other actor-critic algorithms, e.g., TD3 (Fujimoto et al., , 2018).
Since running OK-LS (Alg. 2) until no more corner weights are identified (see line 4) may require a large number of iterations, in the experiments we ran OK-LS for $5$ iterations, which we found to be enough for learning a well-performing meta-policy $\omega$.
To generate sets of test tasks $\mathcal{W}^{\prime}\subset\mathcal{W}$ given different values of $d$, we employed the method introduced by Takagi et al., (2020) available on pymoo (Blank and Deb, , 2020), which produces uniformly-spaced weight vectors in the simplex, $\mathcal{W}$.
E.1 Corner Weights.
Below, we define the concept of corner weights used by OKB (Alg. 1) and OK-LS (Alg. 2), as defined in previous works (Roijers, , 2016; Alegre et al., , 2022).
Definition E.1.
Let $\Psi=\{\bm{\mathbf{\psi}}^{\pi_{i}}\}_{i=1}^{n}$ be a set of SF vectors of $n$ policies. Corner weights are the weights contained in the vertices of a polyhedron, $P$, defined as:
$P=\{\bm{\mathbf{x}}\in\mathbb{R}^{d+1}\ |\ \bm{\mathbf{V}}^{+}\bm{\mathbf{x}}\leq\bm{\mathbf{0}},\textstyle\sum_{i}w_{i}=1,w_{i}\geq 0\},$ (22)
where $\bm{\mathbf{V}}^{+}$ is a matrix whose rows store the elements of $\Psi$ and is augmented by a column vector of $-1$’s. Each vector $\bm{\mathbf{x}}{=}(w_{1},...,w_{d},v_{\bm{\mathbf{w}}})$ in $P$ is composed of a weight vector and its scalarized value.
To compute corner weights, as in Def. E.1, we used pycddlib (https://github.com/mcmtroffaes/pycddlib/) implementation of the Double Description Method (Motzkin et al., , 1953) to efficiently enumerate the vertices of the polyhedron $P$.
E.2 Domains
In this section, we describe in detail the domains used in the experiments, which are shown in Fig. 1.
Minecart domain.
[…]
$\phi_{3}(s,a,s^{\prime})$ $=-0.005-0.025\mathbbm{1}\{a=\mathrm{ACCELERATE}\}-0.05\mathbbm{1}\{a=\mathrm{MiNE}\}.$
We used $\gamma=0.98$ in this domain.
FetchPickAndPlace domain.
We extended the FetchPickAndPlace domain (Plappert et al., , 2018), which consists of a fetch robotic arm that must grab a block on the table with its gripper and move the block to a given target position on top of the table (shown in the middle of Fig. 1). Our implementation of this domain is an adaptation of the one available in Gymnasium-Robotics (de Lazcano et al., , 2023). We note that the state space of this domain, $\mathcal{S}\subset\mathbb{R}^{25}$, is high-dimensional. The action space consists of discretized Manhattan-style movements for the three movement axes, i.e., $\{1:[1.0,0.0,0.0],2:[-1.0,0.0,0.0],3:[0.0,1.0,0.0],4:[0.0,-1.0,0.0],5:[0.0,0.0,1.0],6:[0.0,0.0,-1.0]\}$, and two actions for opening and closing the gripper, totaling $8$ actions. The reward features $\bm{\mathbf{\phi}}(s,a,s^{\prime})\in\mathbb{R}^{d}$ correspond to the negative Euclidean distances between the square block and the $d$ target locations (shown in red in Fig. 1). We used $\gamma=0.95$ in this domain.
Item Collection domain.
This domain consists of a $10\times 10$ grid world, in which an agent (depicted as a yellow circle in the rightmost panel of Fig. 1) moves along the 4 directions, $\mathcal{A}=\{\mathrm{UP},\mathrm{DOWN},\mathrm{LEFT},\mathrm{RIGHT}\}$, and must collect items of two different types (denoted by red triangles and green diamonds, respectively). The reward features $\bm{\mathbf{\phi}}(s,a,s^{\prime})\in\mathbb{R}^{2}$ are indicator functions indicating whether the agent collected one of the items in state $s^{\prime}$. In each episode, $5$ items of each type are randomly placed in the grid. The agent perceives its observation as a $10\times 10$ image with $2$ channels (one per item), which is flattened into a $200$-dimensional vector. We make the observations and dynamics toroidal—that is, the grid “wraps around” connecting cells on opposite edges—similarly as done by Barreto et al., (2019). We used $\gamma=0.95$ in this domain.
Highway domain.
[…]
$\phi_{3}(s,a,s^{\prime})$ $=-1\text{ if }a=\mathrm{TURN\_LEFT}\text{ or }a=\mathrm{TURN\_RIGHT}\text{ else }0.$
All three reward features above are zeroed if the agent is off the road and penalized by $-10$ if the agent crashes into another vehicle. We used $\gamma=0.99$ in this domain.